OpenStudy (anonymous):

factor x^2-8x+16-64y^2

OpenStudy (anonymous):

=(x^2 - 64y^2) - 8x+ 16 = (x - 8y - 4)(x + 8y -4)

OpenStudy (anonymous):

this was a tricky one : x^2 - 64y^2 is the difference of two squares so i used the identity a^2 - b^2 = (a+b)(a-b) which gives (x-8y)(x+8y) we then have to accommodate the - 8x + 16 -4 in each bracket does the does the trick

OpenStudy (anonymous):

i had to do this one on paper before i was confident to type it in

OpenStudy (anonymous):

if you expand the brackets you'll see that its correct